# Explaining Humour Style Classifications: An XAI Approach to Understanding Computational Humour Analysis

### v1

@@ 5 Conclusion and Future Work
This study introduces a comprehensive XAI framework for understanding humour style classification, revealing significant patterns in how linguistic mechanisms and emotional features interact across humour styles. Social forms of humour, particularly affiliative and aggressive styles, demonstrate higher rates of wordplay and semantic conflicts. The emotional signatures vary markedly between styles, with self-enhancing humour showing consistently positive polarity (0.227) and high confidence (0.889), while aggressive humour tends toward negative polarity (-0.039). Our correlation analysis reveals strong relationships between linguistic mechanisms, notably between semantic conflicts and rhyme (r = 0.95), indicating how multiple features work together to create humorous effects. The classification challenges highlight important considerations for future development. The model shows varying confidence levels across styles, particularly struggling with affiliative humour (lowest average confidence: 0.748). These challenges often stem from emotional ambiguity and context misinterpretation, especially when multiple humour mechanisms interact. Self-reference and target identification emerged as critical factors in classification decisions, suggesting their importance for future model development. The practical implications of our findings extend beyond theoretical understanding. Our framework provides actionable insights for researchers studying humour’s role in psychological well-being and social communication, while offering a valuable template for applying XAI to other subjective classification tasks in digital humanities. The framework’s ability to explain model decisions enhances transparency and trust in automated humour analysis, supporting applications in mental health assessment, content moderation, and digital humanities research. Despite the valuable insights provided by this study, several important limitations must be acknowledged. Our analysis focuses on a single model configuration (ALI+XGBoost) with a dataset of 1,463 English-language instances, which inherently limits the generalisability of our findings. The feature detection components faced significant technical constraints, particularly in terms of dictionary coverage and algorithmic accuracy. The CMU Pronouncing Dictionary covered 90.75% of unique words in our dataset for rhyme detection, while WordNet synset coverage reached 88.50% of the vocabulary, leaving notable gaps in handling slang and neologisms. The alliteration detector showed particular limitations, such as considering only the first phoneme, leading to false positives like grouping words beginning with different sounds (e.g., "when"/"one"); counting repeated words as alliteration, and grouping words regardless of their proximity in the text. This contributed to a moderate correlation (r = 0.21) with homonym detection, which is partially due to overlapping feature detection rather than true linguistic relationships. A significant methodological constraint lies in our reliance on pre-trained models for sarcasm (60.7% accuracy), sentiment (93.2% accuracy), and emotion detection (93.8% accuracy), not specifically trained on humorous content. Humorous text contains unique linguistic patterns, irony, and complex emotional layers that may not be well-captured by these general-purpose models. The binary classification of features, such as sarcasm detection, oversimplifies complex linguistic phenomena that exist on a spectrum. Additionally, automated sentiment and emotion detection may not fully capture the subtle, often contradictory emotional cues present in different humour styles, like simultaneous positive and negative emotions in self-deprecating humour. Several analytical constraints affect our findings. The study does not address temporal or contextual variations in humour interpretation, and the analysis of misclassifications may not capture all possible error patterns due to the limited sample size. Future research should expand to multiple models and larger, more diverse datasets, incorporating cross-cultural perspectives and multilingual analysis. Enhanced feature detection algorithms are needed, particularly for complex linguistic features like multi-word puns and contextual wordplay, along with better handling of informal language. The development of specialised models trained specifically on humorous content would improve the accuracy of sentiment and emotion detection in comedic contexts. Additionally, more refined methods for capturing contextual and cultural factors would enhance the framework’s applicability across different contexts, while integration of temporal analysis would provide deeper insights into how humour interpretation varies over time and across different social settings. These improvements would strengthen the framework’s reliability and broaden its applicability in real-world scenarios.
Conspectus librorum
  • Abulaish and Kamal (2018) Muhammad Abulaish and Ashraf Kamal. Self-Deprecating Sarcasm Detection: An Amalgamation of Rule-Based and Machine Learning Approach. In Proceedings - 2018 IEEE/WIC/ACM International Conference on Web Intelligence, WI 2018, pages 574–579. Institute of Electrical and Electronics Engineers Inc., 1 2018. ISBN 9781538673256. 10.1109/WI.2018.00-35.
  • Ahmed et al. (2022) Usman Ahmed, Rutvij H. Jhaveri, Gautam Srivastava, and Jerry Chun-Wei Lin. Explainable Deep Attention Active Learning for Sentimental Analytics of Mental Disorder. ACM Transactions on Asian and Low-Resource Language Information Processing, 7 2022. ISSN 2375-4699. 10.1145/3551890.
  • Amjad and Dasti (2022) Arooba Amjad and Rabia Dasti. Humor styles, emotion regulation and subjective well-being in young adults. Current Psychology, 41(9):6326–6335, 9 2022. ISSN 19364733. 10.1007/s12144-020-01127-y.
  • Anderson and Di Tunnariello (2016) Whitney Anderson and Nancy Di Tunnariello. Aggressive humor as a negative relational maintenance behavior during times of conflict. Qualitative Report, 21(8):1513–1530, 8 2016. ISSN 21603715. 10.46743/2160-3715/2016.2149.
  • Annamoradnejad and Zoghi (2020) Issa Annamoradnejad and Gohar Zoghi. ColBERT: Using BERT Sentence Embedding in Parallel Neural Networks for Computational Humor. 4 2020. URL http://arxiv.org/abs/2004.12765.
  • Chen and Martin (2007) Guo Hai Chen and Rod A. Martin. A comparison of humor styles, coping humor, and mental health between Chinese and Canadian university students. Humor, 20(3):215–234, 8 2007. ISSN 09331719. 10.1515/HUMOR.2007.011.
  • Chen et al. (2024) Yuyan Chen, Yichen Yuan, Panjun Liu, Dayiheng Liu, Qinghao Guan, Mengfei Guo, Haiming Peng, Bang Liu, Zhixu Li, and Yanghua Xiao. Talk Funny! A Large-Scale Humor Response Dataset with Chain-of-Humor Interpretation. In The Thirty-Eighth AAAI Conference on Artificial Intelligence (AAAI-24), pages 17826–17834, 2024. URL https://openai.com/blog/chatgpt.
  • Chowdhury et al. (2021) Kounteyo Roy Chowdhury, Arpan Sil, and Sharvari Rahul Shukla. Explaining a Black-Box Sentiment Analysis Model with Local Interpretable Model Diagnostics Explanation (LIME). In Communications in Computer and Information Science, volume 1440 CCIS, pages 90–101. Springer Science and Business Media Deutschland GmbH, 2021. ISBN 9783030814618. 10.1007/978-3-030-81462-5_9.
  • Cortinas-Lorenzo and Lacey (2024) Karina Cortinas-Lorenzo and Gerard Lacey. Toward Explainable Affective Computing: A Review. IEEE Transactions on Neural Networks and Learning Systems, 2024. ISSN 21622388. 10.1109/TNNLS.2023.3270027.
  • De Marez et al. (2024) Victor De Marez, Thomas Winters, and Ayla Rigouts Terryn. THInC: A Theory-Driven Framework for Computational Humor Detection. In InternationalWorkshop on Artificial Intelligence and Creativity, Santiago de Compostela,, 9 2024. URL http://arxiv.org/abs/2409.01232.
  • Edalat (2023) Abbas Edalat. Self-initiated humour protocols: An algorithmic approach for learning to laugh. PsyArXiv, 5:1–14, 11 2023. https://doi.org/10.31234/osf.io/w9cvx.
  • Edalat et al. (2024) Abbas Edalat, Ruoyu Hu, Zeena Patel, Neophytos Polydorou, Frank Ryan, and Dasha Nicholls. Self-Initiated Humour Protocol: A pilot study with an AI agent. PsyArXiv, 2024. 10.31234/osf.io/f32xq.
  • Hartmann et al. (2023) Jochen Hartmann, Mark Heitmann, Christian Siebert, and Christina Schamp. More than a Feeling: Accuracy and Application of Sentiment Analysis. International Journal of Research in Marketing, 40(1):75–87, 3 2023. ISSN 01678116. 10.1016/j.ijresmar.2022.05.005.
  • Kamal and Abulaish (2020) Ashraf Kamal and Muhammad Abulaish. Self-deprecating Humor Detection: A Machine Learning Approach. In Le-Minh Nguyen, Xuan-Hieu Phan, Kôiti Hasida, and Satoshi Tojo, editors, Computer Lingustics, volume 1215 of Communications in Computer and Information Science, pages 483–484. Springer Singapore, Singapore, 2020. ISBN 978-981-15-6167-2. 10.1007/978-981-15-6168-9. URL http://link.springer.com/10.1007/978-981-15-6168-9.
  • Kazienko et al. (2023) Przemysław Kazienko, Julita Bielaniewicz, Marcin Gruza, Kamil Kanclerz, Konrad Karanowski, Piotr Miłkowski, and Jan Kocoń. Human-centered neural reasoning for subjective content processing: Hate speech, emotions, and humor. Information Fusion, 94:43–65, 6 2023. ISSN 15662535. 10.1016/j.inffus.2023.01.010.
  • Kenneth et al. (2024) Mary Ogbuka Kenneth, Foaad Khosmood, and Abbas Edalat. Systematic Literature Review: Computational Approaches for Humour Style Classification. Technical report, 2024. URL https://www.semanticscholar.
  • Kenneth Ogbuka et al. (2024) Mary Kenneth Ogbuka, Foaad Khosmood, and Abbas Edalat. A Two-Model Approach for Humour Style Recognition. In Proceedings of the 4th International Conference on Natural Language Processing for Digital Humanities, pages 259–274, Miami, 11 2024. Association for Computational Linguistics. URL https://aclanthology.org/2024.nlp4dh-1.25/.
  • Kuiper et al. (2016) Nicholas Kuiper, Gillian Kirsh, and Nadia Maiolino. Identity and Intimacy Development, Humor Styles, and Psychological Well-Being. Identity, 16(2):115–125, 4 2016. ISSN 1532706X. 10.1080/15283488.2016.1159964.
  • Lundberg and Lee (2017) Scott Lundberg and Su-In Lee. A Unified Approach to Interpreting Model Predictions. In 31st Conference on Neural Information Processing Systems, Long Beach, 5 2017. URL http://arxiv.org/abs/1705.07874.
  • Lyu et al. (2024) Qing Lyu, Marianna Apidianaki, and Chris Callison-Burch. Towards Faithful Model Explanation in NLP: A Survey. Computational Linguistics, pages 657–723, 2024. 10.1162/coli. URL https://doi.org/10.1162/coli.
  • Mahajan and Zaveri (2024) Rutal Mahajan and Mukesh Zaveri. An automatic humor identification model with novel features from Berger’s typology and ensemble models. Decision Analytics Journal, 11, 6 2024. ISSN 27726622. 10.1016/j.dajour.2024.100450.
  • Mann and Mikulandric (2024) Rowan Mann and Tomislav Mikulandric. CLEF 2024 JOKER Tasks 1-3: Humour identification and classification. In Conference and Labs of the Evaluation Forum, Grenoble, 9 2024.
  • Martin and Ford (2018) Rod A Martin and Thomas Ford. The psychology of humor: An integrative approach. Academic Press, 2nd edition, 2018.
  • Martin et al. (2003) Rod A Martin, Patricia Puhlik-Doris, Gwen Larsen, Jeanette Gray, and Kelly Weir. Individual differences in uses of humor and their relation to psychological well-being: Development of the Humor Styles Questionnaire. Journal of Research in Personality, 37:48–75, 2003. URL www.elsevier.com/locate/jrp.
  • Ortega-Bueno et al. (2022) Reynier Ortega-Bueno, Paolo Rosso, and José E. Medina Pagola. Multi-view informed attention-based model for Irony and Satire detection in Spanish variants. Knowledge-Based Systems, 235, 1 2022. ISSN 09507051. 10.1016/j.knosys.2021.107597.
  • Pérez-Landa et al. (2021) Gabriel Ichcanziho Pérez-Landa, Octavio Loyola-González, and Miguel Angel Medina-Pérez. An explainable artificial intelligence model for detecting xenophobic tweets. Applied Sciences (Switzerland), 11(22), 11 2021. ISSN 20763417. 10.3390/app112210801.
  • Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. "Why Should I Trust You?" Explaining the Predictions of Any Classifier. In Proceedings of NAACL-HLT, pages 97–101, San Diego, 6 2016. Association for Computational Linguistics. URL https://github.
  • Sari (2016) Serkan Volkan Sari. Was it just joke? Cyberbullying perpetrations and their styles of humor. Computers in Human Behavior, 54:555–559, 1 2016. ISSN 07475632. 10.1016/j.chb.2015.08.053.
  • Sen (2012) Anindya Sen. Humour Analysis and Qualitative Research. Technical report, University of Surrey, 2012.
  • Zhang et al. (2017) Donghai Zhang, Wei Song, Lizhen Liu, Chao Du, and Xinlei Zhao. Investigations in automatic humor recognition. In Proceedings - 2017 10th International Symposium on Computational Intelligence and Design, ISCID 2017, volume 1 2018-January, pages 272–275. Institute of Electrical and Electronics Engineers Inc., 7 2017. ISBN 9781538636749. 10.1109/ISCID.2017.160.
  • Zhu et al. (2022) Hong Zhu, Yilin Ou, and Zimeng Zhu. Aggressive humor style and cyberbullying perpetration: Normative tolerance and moral disengagement perspective. Frontiers in Psychology, 13, 12 2022. ISSN 16641078. 10.3389/fpsyg.2022.1095318.

### v2

@@ 5 Conclusion and Future Work
This study introduces a comprehensive XAI framework for understanding humour style classification, revealing significant patterns in how linguistic mechanisms and emotional features interact across humour styles. Social forms of humour, particularly affiliative and aggressive styles, demonstrate higher rates of wordplay and semantic conflicts. The emotional signatures vary markedly between styles, with self-enhancing humour showing consistently positive polarity (0.227) and high confidence (0.889), while aggressive humour tends toward negative polarity (-0.039). Our correlation analyses reveal complex relationships between linguistic mechanisms, particularly in the relationship between semantic conflicts and rhyme, where Pearson correlation (r = 0.95) suggests a strong linear relationship, but the much lower Spearman correlation (r = 0.32) indicates this relationship is heavily influenced by outliers rather than being typical across most instances.
The classification challenges highlight important considerations for future development. The model shows varying confidence levels across styles, particularly struggling with affiliative humour (lowest average confidence: 0.748). These challenges often stem from emotional ambiguity and context misinterpretation, especially when multiple humour mechanisms interact. Self-reference and target identification emerged as critical factors in classification decisions, suggesting their importance for future model development.
The practical implications of our findings extend beyond theoretical understanding. Our framework provides actionable insights for researchers studying humour’s role in psychological well-being and social communication, while offering a valuable template for applying XAI to other subjective classification tasks in digital humanities. The framework’s ability to explain model decisions enhances transparency and trust in automated humour analysis, supporting applications in mental health assessment, content moderation, and digital humanities research.
Despite the valuable insights provided by this study, several important limitations must be acknowledged. Our analysis focuses on a single model configuration (ALI+XGBoost) with a dataset of 1,463 English-language instances, which inherently limits the generalisability of our findings. The feature detection components faced significant technical constraints, particularly in terms of dictionary coverage and algorithmic accuracy. The CMU Pronouncing Dictionary covered 90.75% of unique words in our dataset for rhyme detection, while WordNet synset coverage reached 88.50% of the vocabulary, leaving significant gaps in handling slang and neologisms. The alliteration detector showed particular limitations, such as considering only the first phoneme, leading to false positives like grouping words beginning with different sounds (e.g., “when/one”), counting repeated words as alliteration, and grouping words regardless of their proximity in the text. This contributed to moderate correlations with homonym detection (Pearson: 0.21, Spearman: 0.27), which is partially due to overlapping feature detection rather than true linguistic relationships.
A significant methodological constraint lies in our reliance on pre-trained models for sarcasm (60.7% accuracy), sentiment (93.2% accuracy), and emotion detection (93.8% accuracy), not specifically trained on humorous content. Humorous text contains unique linguistic patterns, irony, and complex emotional layers that may not be well-captured by these general-purpose models. The binary classification of features, such as sarcasm detection, oversimplifies complex linguistic phenomena that exist on a spectrum. Additionally, automated sentiment and emotion detection may not fully capture the subtle, often contradictory emotional cues present in different humour styles, like simultaneous positive and negative emotions in self-deprecating humour.
Several analytical constraints affect our findings. The study does not address temporal or contextual variations in humour interpretation, and the analysis of misclassifications may not capture all possible error patterns due to the limited sample size. Future research should expand to multiple models and larger, more diverse datasets, incorporating cross-cultural perspectives and multilingual analysis. Enhanced feature detection algorithms are needed, particularly for complex linguistic features like multi-word puns and contextual wordplay, along with better handling of informal language. The development of specialised models trained specifically on humorous content would improve the accuracy of sentiment and emotion detection in comedic contexts. Additionally, more refined methods for capturing contextual and cultural factors would enhance the framework’s applicability across different contexts, while integration of temporal analysis would provide deeper insights into how humour interpretation varies over time and across different social settings. These improvements would strengthen the framework’s reliability and broaden its applicability in real-world scenarios.
6 Acknowledgement
This research was supported by the Petroleum Technology Development Fund (PTDF) of Nigeria.
Appendix A Summary of Related works
Table 5 summarises the key works discussed in section 2 (Related works), highlighting the authors’ specific tasks, datasets used, features extracted, classification approaches, XAI methods employed, and their achieved results. This overview demonstrates both the progress made in computational humour analysis and explainability techniques while highlighting the gaps our work addresses.
Pérez-Landa et al. [2021] Xenophobia
Experts Xenophobia
Database (EXD) (10,057 tweets)
Pitropakis Xenophobia Database
(PXD)(5814 tweets)
TF-IDF, Bag of words (BOW),
Word 2 vector (W2V),
Keywords, Sentiment,
Emotion, and Syntactic
C45, KNN, Rusboost,
UnderBagging,
Feature engineering
(keywords, sentiment,
emotions and syntactic
F1: 76.8% (EXD)
AUC: 0.864 (EXD)
F1: 73.4% (PXD)
AUC: 0.794 (PXD)
Chowdhury et al. [2021] Sentiment Twitter API (2-years period)
W2V, Sentiment Specific Word
Embedding (SSWE), GloVe,
FastText embeddings
Bi-directional LSTM LIME
ACC: 72%
F1: 72%
Ahmed et al. [2022] Sentiment Patient-authored text (15,044)
Fuzzy logic rules, GloVe,
Cosine Similarity, Keywords
Bi-directional LSTM
Attention with fuzzy
classification
Fuzzy logic rules F1: 89%
Ortega-Bueno et al. [2022] Irony, Satire
IroSvA’19 shared task (3000)
Barbieri’15-es (10,000)
Salas’17-mx (10,888)
HAHA’19 shared task (6000)
BERT embedding,
Multilingual Universal
Sentence encoding,
semantics, affective, incongruity,
stylistic and structural
Multiview Attention LSTM
Feature engineering
(semantics, affective,
incongruity, stylistic
and structural)
F1: 70.4% (IroSva’19)
F1: 95.7% (Barbieri’15))
F1: 96% (Salas’17)
F1: 80.6% (HAHA’19)
Zhang et al. [2017] Humour 16000 One-Liner (32,002)
Contextual knowledge,
Subjectivity, affective polarity,
ambiguity, incongruity
Feature engineering
(Contextual knowledge,
Subjectivity, polarity,
ambiguity, incongruity)
F1: 85%
Mann and Mikulandric [2024] Humour, Pun CLEF 2024 JOKER Tasks TF-IDF
Logistic regression,
Naive Bayes (NB),
SVM, MarianMTModel
F1: 83% (Pun)
F1: 61% (Humour)
De Marez et al. [2024] Humour SemEval 2021 Task 7
Polarity, emotions, offense,
subjectivity, hate, stance,
ambiguity, adult language
GA2M
Feature engineering
(Polarity, emotions,
offense, subjectivity,
hate, stance, ambiguity,
adult language)
F1: 85%
Mahajan and Zaveri [2024] Humour Yelp reviews Emotive, incongruity, intensity
Stacking-based ensemble,
NB, SVM, MLP,
majority-vote ensemble
Feature engineering
(Berger’s typology-based
incongruity features)
F1: 72.57%
Chen et al. [2024] Humour Chain-of-humour dataset word embeddings
Pre-trained language
models (PLMs)
Chain-of-thought F1: 87%
Abulaish and Kamal [2018] Sarcasm Twitter API (7 weeks tweet)
Self-referencing, hyperbolic,
part-of-speech
Decision tree,
Naive Bayes,
Bagging
F1: 94%
Kamal and Abulaish [2020] Humour
Pun of the Day (4,826),
1600 One-Liner (32.002),
Twitter API (20,000)
Self-deprecating pattern,
exaggeration, word embedding,
ambiguity, interpersonal effect,
phonetic style.
Random forest
F1: 62% - 87%
across datasets
Kenneth Ogbuka et al. [2024] Humour Styles Humour styles dataset Sentence embedding
NB, Randon forest,
XGBoost, DistilBERT
F1: 78.6%
Table 5: Summary of Related Works